A 14.5 L sample of gas at 1.6 atm and 26.00 C is heated to 97.0 0 C and transferred to a container where its pressure is 1.00 atm. What is the volume of the container?

1
Expert's answer
2021-04-01T06:01:34-0400

"\frac {P1V1}{T1}=\frac{P2V2}{T2}"


"V2" is the volume of container


"V2 = \frac{P1V1T2}{T1P2}"


T1 = 273+26=299K

T2= 273+97=370k


"V2= \frac{1.6×14.5×370}{299×1.00}=28.7L"